an object move with a uniform speed v1 for a long time t  and move with a uniform speed v2 futher time t find the average speed

Grade:9

1 Answers

noogler
8 years ago
L1=v1t , L2=v2t
Vavg=tot.dis./tot. time = L1+L2/t+t = (v1+v2)t/2t = v1+v2/2
avg. speed=v1+v2/2
 
APPROVE IF USEFULL!!!!!!!!!!!!!